# Quantum-Enhanced Processing with Tensor-Network Frontends for Privacy-Aware Federated Medical Diagnosis

### 02 Apr 2026 v1

@@ I Introduction
The rapid digitalization of healthcare, finance, and public infrastructure has led to a sustained increase in distributed, privacy-sensitive data. In medical domains in particular, electronic health records, imaging repositories, and wearable sensor streams are inherently decentralized across hospitals and institutions [27, 28]. At the same time, regulatory frameworks such as the General Data Protection Regulation (GDPR) [11], the Health Insurance Portability and Accountability Act (HIPAA) [33], and national medical data protection laws including Japan’s Act on the Protection of Personal Information (APPI) [14] and the Next-Generation Medical Infrastructure Act [15] impose strict restrictions on data centralization and cross-institutional sharing. Recent policy initiatives such as the European Health Data Space (EHDS) [10] further reinforce the need for controlled secondary use of medical data under strong governance requirements. As a result, modern medical AI faces a central tension: high-performance learning benefits from access to broad and heterogeneous data, whereas direct data pooling is increasingly constrained.
Federated learning (FL) has therefore emerged as an important paradigm for privacy-sensitive collaborative learning because it allows multiple institutions to participate in joint model training without transferring raw data to a central repository [22, 16]. This setting is especially relevant in medicine, where data locality is often a regulatory and operational requirement rather than a design preference [27]. However, data locality alone does not fully resolve the privacy problem. Even when raw inputs remain on site, gradients, intermediate representations, and model updates can still leak sensitive information through inversion and inference attacks [37, 29]. Because such attacks rely on access to individual client updates, restricting visibility to aggregated information becomes an important design principle.
Secure aggregation based on multi-party computation (MPC) mitigates this risk by ensuring that only aggregated client contributions become available [4]. In MPC, a value is decomposed into secret shares distributed across separate non-colluding computation nodes, which jointly execute a protocol on protected inputs. The final output is reconstructed only from the shares, while no party learns more than what is implied by the released result. MPC is therefore a natural choice for protecting client contributions. Moreover, because this work investigates a constructive use of quantum computation, it is also necessary to consider the possibility that quantum computation could be exploited by a malicious adversary to compromise privacy. The privacy mechanism should therefore remain secure against quantum-computer-aided attackers. Secret-sharing-based MPC, including the variant considered in this work, provides unconditional security. Accordingly, the MPC model adopted here is secure against any computational adversary, including quantum-computational ones.
While privacy in the present framework is provided by MPC-secured aggregation, an additional security question arises if the post-aggregation Quantum-Enhanced Processor (QEP) is delegated to an external quantum server. Blind quantum computing (BQC), including enclave-based approaches such as QEnclave, offers a possible way to hide such delegated quantum computations from the server [5, 13, 21], but this lies outside the present scope.
Although MPC strengthens privacy protection, it also introduces a nontrivial systems trade-off. Expanding the protected portion of the pipeline, or adopting stronger adversarial assumptions, generally increases communication and execution overhead. This raises a fundamental architectural question for regulated federated learning: how should privacy strength, representation dimension, and downstream inference quality be balanced within a single system? The question becomes even more consequential when quantum enhancement is introduced, because any quantum module must operate under the constraints imposed by secure aggregation while remaining compatible with the structure of the aggregated latent representation.
In parallel, hybrid quantum machine learning has attracted attention as a framework for structured nonlinear feature transformation. Distributed quantum neural network models [25], communication-efficient quantum learning algorithms [30], and broader studies of distributed quantum computing and simulation [2, 12] suggest that quantum-enhanced learning should be evaluated together with systems-level scalability considerations. Hybrid tensor-network and quantum-classical models [6, 20, 32, 7, 19, 31] further suggest that structured classical representations can serve as effective interfaces to downstream quantum feature-processing modules. In privacy-preserving quantum federated settings, Li et al. proposed quantum-native federated protocols based on gradient hiding and private inner-product estimation [18]. Nevertheless, prior studies have typically emphasized either protocol design or hybrid model design in isolation, and they do not sufficiently characterize how a post-aggregation quantum refinement module behaves when embedded in a privacy-aware federated learning pipeline that simultaneously includes representation compression and secure aggregation.
A further challenge arises from current hardware limitations. Available quantum processors provide only a limited number of usable qubits, so directly processing large medical-image inputs within a quantum module is generally unrealistic. In practice, a useful hybrid design must make effective use of a small quantum system rather than assume access to large fault-tolerant devices. This requires a representation strategy that compresses the original input before quantum processing while preserving task-relevant information.
Tensor networks (TNs) are a natural candidate for this role. Matrix Product States (MPS) were introduced as expressive supervised learning models in [17], and subsequent work explored spectral tensor-train parameterizations [24], residual tensor-train architectures [8], and MPS pre-training strategies for quantum learning [9]. Tensor networks have also been applied to generative modeling [23], high-energy physics event reconstruction [1], and sequential state learning [36]. Recent work has additionally examined TNs in federated and distributed settings, including coupled tensor-train decomposition for federated learning [35] and hierarchical tensor-network-based federated frameworks [3]. Collectively, these studies establish TNs as compact alternatives to dense neural architectures with explicit structural inductive bias. In the present setting, their role is twofold. First, they compress large client-side inputs into low-dimensional latent representations, making post-aggregation quantum processing feasible with a small number of qubits. Second, because the protected representation itself becomes smaller, they reduce the communication overhead incurred by MPC on that representation.
@@
Despite these promising ingredients, prior work has not sufficiently integrated them within a single privacy-aware federated learning pipeline. Two questions remain particularly unresolved: whether compressed latent representations can simultaneously improve the practicality of MPC and support effective small-qubit quantum processing, and whether the benefit of post-aggregation quantum refinement depends on the tensor-network architecture that produces the latent representation.
These issues motivate the present study. The first unresolved problem is practical: MPC can improve privacy in federated learning, but its communication overhead becomes burdensome when high-dimensional representations must be protected. The second unresolved problem is architectural: if only a small number of qubits is available, then the utility of a quantum module depends critically on how the input is compressed before quantum processing. The third unresolved problem is empirical: it remains unclear whether the effect of post-aggregation quantum enhancement is architecture-independent or instead determined by its interaction with the frontend representation.
@@
To address these issues, we propose an end-to-end privacy-aware federated learning framework for medical image classification that combines client-side tensor-network encoding, MPC-secured aggregation of latent representations, and post-aggregation quantum refinement by a Quantum-Enhanced Processor (QEP). We compare three tensor-network frontends—MPS, TTN, and MERA—and evaluate how the effect of the QEP depends on the latent structure supplied by the frontend. The resulting design is intended to support small-qubit quantum processing while reducing the communication overhead associated with secure aggregation.
Experiments are conducted on PneumoniaMNIST [34], a compact biomedical benchmark suitable for controlled evaluation of privacy-aware federated learning pipelines. Our analysis yields two main findings. First, the effect of the QEP depends on its pairing with the tensor-network frontend rather than on the quantum module alone. In particular, the combination of TTN and QEP shows the most favorable overall profile in the present setting, with the clearest positive tendency under quantum enhancement. By contrast, MPS, TTN, and MERA occupy distinct operating regimes in their internal behavior, indicating that the impact of quantum enhancement is architecture-dependent. Second, the MPC benchmark shows that communication cost is governed primarily by the dimension of the protected representation. This means that tensor-network compression not only makes small-qubit quantum processing feasible, but also improves the practicality of secure deployment by reducing the communication overhead associated with MPC.
The central premise of this study is therefore that representation compression, secure aggregation, and quantum refinement should not be treated as independent design variables. Compressing the input before aggregation changes the cost of privacy protection; the resulting latent structure conditions the behavior of the quantum module; and the practical value of the overall pipeline depends on balancing predictive performance, quantum operating behavior, and secure-computation cost. From this perspective, the main contribution of this work is not a standalone quantum algorithm, but a co-designed system showing how tensor-network compression can support small-qubit quantum enhancement in privacy-aware federated learning.
To summarize, the main contributions of this work are as follows:
@@
  • We present an end-to-end privacy-aware hybrid learning framework that combines tensor-network latent encoding, MPC-secured aggregation, and post-aggregation quantum refinement for federated medical image classification.
  • We show that tensor-network compression plays a dual role: it supports small-qubit quantum processing on compressed latent features and simultaneously reduces the communication overhead of MPC-secured aggregation.
  • We demonstrate that the effect of the Quantum-Enhanced Processor is architecture-dependent and that, in the present setting, the TTN+QEP combination provides the most favorable overall profile, including improved performance relative to its classical counterpart.
@@ II Methodology
Figure 1: End-to-end overview of the proposed framework. Client-side tensor-network embedding compresses each local input into a latent representation, MPC-secured aggregation and gating produce a protected global latent feature, and the Quantum-Enhanced Processor (QEP) performs post-aggregation quantum refinement before final decision making. Experiment 1 evaluates the inference performance of this pipeline across different tensor-network frontends, while Experiment 2 analyzes the communication cost of the MPC-secured aggregation component through an abstract protected-aggregation model.
Fig. 1 shows the proposed privacy-aware federated learning framework. The architecture consists of three coupled stages: client-side tensor-network encoding, MPC-secured aggregation of latent representations, and post-aggregation refinement by a Quantum-Enhanced Processor (QEP). The framework is formulated as an integrated TN+MPC+QEP pipeline so that representational compression, protected aggregation, and quantum refinement can be analyzed within a unified setting.
At the client side, each local input is transformed into a compact latent representation by a tensor-network frontend. These latent features are then combined through an MPC-secured aggregation layer that produces a protected global representation, which is subsequently processed by the QEP before final classification. Experiment 1 evaluates the end-to-end predictive behavior of this system under different tensor-network frontends, whereas Experiment 2 isolates the MPC-secured aggregation stage and analyzes its communication cost under progressively stronger security settings.
@@ II-A Client-Side Tensor-Network Embedding
For the TTN and MERA branches, the input is first reorganized into non-overlapping image patches rather than being directly reshaped as a one-dimensional vector sequence. In particular, the $28\times 28$ image is partitioned into $16$ patches of size $7\times 7$, and each patch is processed by a shared patch-wise stem network:
@@
Accordingly, the client-side tensor-network block serves two roles. First, it compresses the original image into a lower-dimensional latent representation before aggregation, reducing the dimensional burden of the subsequent server-side pipeline. Second, it determines the latent geometry on which the Quantum-Enhanced Processor (QEP) later operates. In this sense, the TN frontend is not merely a compression stage, but the structural interface that conditions downstream quantum refinement.
@@ II-A1 MPS
In the implementation, the site tensors are parameterized in a QR-projected left-canonical form, so that the matricized tensor approximately satisfies
$(A^{(k)})^{\dagger}A^{(k)}\approx I.$ (10)
This constraint improves numerical stability during sequential contraction. In addition, the intermediate hidden state is normalized at each step to control norm growth or collapse.
After the final contraction, the resulting complex state is converted to a real-valued feature and projected to the shared latent dimension $d$.
@@ II-A2 TTN
where $\cdot\|\cdot$] denotes vector concatenation. In the implementation, $\mathcal{W}^{(\ell)}$ is parameterized as a complex isometric transformation obtained through QR-based projection, so that
$(\mathcal{W}^{(\ell)})^{\dagger}\mathcal{W}^{(\ell)}\approx I.$ (15)
After each aggregation level, the hidden state is normalized for numerical stability.
After the final contraction, the root state is converted to a real-valued feature and projected to $\mathbb{R}^{d}$.
@@ II-A3 MERA
followed by a split back into two updated local states. In the implementation, these disentanglers are applied in two stages: first on even pairs $(0,1),(2,3),\dots$, and then on odd pairs $(1,2),(3,4),\dots$. This alternating pattern changes the local coupling structure across each level without requiring a cyclic shift.
@@
where $\hat{h}^{(\ell)}$ denotes the disentangled local states. Both the disentangling and coarse-graining maps are implemented as QR-projected complex transformations, and the hidden states are normalized after each coarse-graining step for numerical stability.
After the final level, the root state is converted to a real-valued latent representation and projected to $\mathbb{R}^{d}$. Although the tensor-network frontend is described at the modeling level over real-valued latent features (with complex internal tensor states where appropriate), the discrete fixed-point assumption is introduced only for the MPC benchmark; the effect of explicitly quantizing TN outputs before secure computation is not evaluated in the present study.
@@ II-B1 System Model and Threat Model
We consider $n$ clients $\{C_{i}\}_{i=1}^{n}$ and an aggregation server $S$. Each client holds a private dataset $\mathcal{D}_{i}$ and computes a local representation $f_{i}\in\mathbb{R}^{d}$. For the MPC cost analysis, we additionally introduce a client-specific scalar coefficient $w_{i}\in\mathbb{R}$ to model protected weighted aggregation and normalization, as described in Section II-C.
To protect client data, we use an outsourced 3-party MPC model in which the role of $S$ is distributed across three non-colluding computation nodes $S_{0},S_{1},S_{2}$. We consider both passive and active security. Scenarios 1–3 correspond to passive-security settings, whereas Scenarios 4–6 are their active-security counterparts.
@@ II-B2 Cost Model
In the communication model used in this paper, one secure multiplication costs $3k$ bits, one truncation costs $6k$ bits, and a full fixed-point multiplication therefore costs $9k$ bits. Secure division is modeled with communication cost $3k(k+4\theta+2)$, where $\theta$ denotes the number of refinement iterations. Active security is modeled by doubling the communication cost. Because our focus is communication benchmarking rather than protocol derivation, we use these established costs directly and do not reproduce the full protocol details here.
@@ II-C Secure Aggregation via MPC
The proposed end-to-end framework includes an MPC-secured aggregation stage between client-side tensor-network embedding and post-aggregation quantum refinement. For inference evaluation, this functionality is studied through the system-level architecture shown in Fig. 1. For communication-cost analysis, we abstract the aggregation stage into a protected weighted aggregation form so that its security and communication overhead can be studied independently of low-level implementation details.
@@ II-D Quantum-Enhanced Processor
The central quantum component of the framework is the Quantum-Enhanced Processor (QEP), which refines the aggregated latent representation through a hybrid quantum-classical transformation. Let the aggregated server-side feature be denoted by
@@
The QEP maps $\mathbf{x}_{\mathrm{agg}}$ to a quantum-derived feature vector through a quantum-state embedding followed by observable-based readout, and fuses it with the original classical representation via adaptive residual mixing. In the present implementation, the quantum circuit is not treated as a trainable component. Instead, it is evaluated using statevector simulation and used as a fixed nonlinear feature transformation that deterministically maps the input to quantum-derived statistics.
As a result, gradient-based optimization is not applied to the quantum circuit itself. Instead, only the surrounding classical components—including the encoder, decoder, and fusion modules—are trained to determine how the quantum-derived features are incorporated into the latent representation. This design allows the QEP to function as a quantum feature extractor rather than as a variational quantum model.
From a representational viewpoint, the QEP can be interpreted as a feature-expansion mechanism that embeds a compact latent vector into a quantum Hilbert space and extracts observable-based statistics. In this setting, the effective dimensionality of the quantum feature space is determined by the number of observables used for measurement rather than by the number of circuit parameters. To ensure sufficient expressivity, it is therefore desirable that the number of quantum features $d_{q}$ be comparable to the latent dimension $d$. When using both one-body and two-body Pauli observables, this leads to a scaling relation $d_{q}=\mathcal{O}(N_{q}^{2})$, suggesting that the number of qubits should scale approximately as $N_{q}\sim\sqrt{d}$. This consideration motivates the choice of qubit counts in the present study, and its empirical validity is examined in the qubit-scaling experiments described later.
@@ II-D1 Quantum Encoding and Observable Readout
where $d_{q}$ is determined by the number and type of observables. In practice, this allows the model to extract higher-order correlations and nonlinear features from a compact latent representation without requiring a large quantum register.
@@ II-D2 Classical Decoding and Residual Fusion
In this way, the model can adaptively control the contribution of the quantum-enhanced branch rather than relying on it uniformly across all samples. The QEP is therefore not only a simulated quantum feature map, but also a learned controller over how strongly quantum-derived information influences the final latent representation.
@@ II-D3 Training Objective
Only the surrounding classical components—including the encoder, decoder, bypass pathway, and fusion modules—are optimized during training, while the quantum circuit serves as a fixed observable-based feature transformation.
@@ III Experimental Setup
We evaluated the proposed framework on PneumoniaMNIST, a binary medical image classification benchmark from the MedMNIST collection [34]. The dataset is derived from pediatric chest X-ray images, and the task is to discriminate Normal from Pneumonia. Following the official preprocessing protocol, each image was resized to $28\times 28$ pixels, yielding a single-channel grayscale input. In the implementation, each image was normalized to the range $[0,1]$. For the MPS branch, the image was flattened into a 784-dimensional vector. For the TTN and MERA branches, the same image was additionally reorganized into $16$ non-overlapping patches of size $7\times 7$, followed by a shared patch-wise stem transformation. The standard training, validation, and test splits provided by the dataset interface were used throughout the experiments. Fig. 3 shows representative examples of the resized grayscale chest X-ray images in PneumoniaMNIST, including both Normal and Pneumonia cases.
Figure 3: Example samples from PneumoniaMNIST. The top row shows Normal cases and the bottom row shows Pneumonia cases. All images are resized grayscale chest X-ray images used in the present experiments.
To reflect the distributed setting of the proposed architecture, the data in each split were partitioned across multiple client branches using label-stratified client assignment. Unless otherwise noted, the number of client branches was fixed at 16. Training used a per-client local batch size of 4, yielding up to 64 samples per federated training step when all client branches contributed non-empty batches. This configuration is consistent with the client-side tensor-network embedding pipeline described in Section II-A, where each branch produces a compressed latent representation prior to server-side gating and optional quantum enhancement.
The overall model consisted of three stages: a client-side tensor-network embedding module, a server-side feature-gating module, and a Quantum-Enhanced Processor (QEP), followed by a classical classifier. For the client-side tensor-network interface, we considered three architectures: Matrix Product State (MPS), Tree Tensor Network (TTN), and Multi-scale Entanglement Renormalization Ansatz (MERA). In all cases, the shared latent feature dimension was fixed to 64. The primary goal of the experiments was to characterize how the QEP behaves across these frontend-induced latent structures.
@@
For the QEP, the quantum branch used 16 qubits, circuit depth 2, angle scale 0.5, and refresh interval 1. Quantum expectation values were computed with Qiskit Aer in statevector mode. The circuit consisted of repeated layers of single-qubit $R_{y}$ and $R_{z}$ rotations followed by nearest-neighbor CNOT entangling gates, and the observable set included one-body Pauli terms together with selected two-body correlations. The resulting quantum statistics were classically decoded into the shared latent feature space and adaptively fused with the classical pathway.
We report both standard-threshold and threshold-optimized evaluation. In the standard setting, the Pneumonia posterior was thresholded at 0.5. In the optimized setting, the threshold was selected on the validation split from the ROC curve using Youden’s $J$ statistic and then applied to the test split. Performance was assessed using Accuracy, Precision, Recall, and F1-score for both Normal and Pneumonia classes. For quantum runs, we additionally analyzed internal QEP diagnostics, including the mean fusion coefficient $\alpha$, the standard deviation of the quantum-enhanced branch representation ($q$-std), the mean magnitude of the quantum feature vector, and the cosine similarity between classical and quantum features, in order to characterize the internal operating regime of the QEP.
We also examined the behavior of the QEP under qubit scaling and noise variation. To study scalability with respect to quantum register size, we varied the number of qubits from $4$ to $16$ while keeping the remaining training and evaluation protocol unchanged. To study noise sensitivity, we fixed the QEP to $8$ qubits and compared noiseless simulation, depolarizing noise, thermal noise, and a mixed-noise configuration. Because the present QEP is implemented as a simulation-based observable feature generator, these analyses should be interpreted as controlled studies of the operating range and robustness of the quantum branch rather than as hardware-performance claims.
The comparison in this study is intended to evaluate practically instantiated frontend families rather than a strictly topology-isolated control in which all preprocessing components are forced to be identical. Accordingly, each frontend is implemented in a form suited to its own structural bias, and the comparison is used to examine how different tensor-network representation regimes interact with post-aggregation quantum refinement within the same end-to-end framework.
Based on this common training and evaluation setting, we define two complementary experiments: one for end-to-end predictive analysis and one for MPC communication-cost benchmarking.

### 03 Apr 2026 v2

@@ I Introduction
The rapid digitalization of healthcare has increased the volume of distributed, privacy-sensitive medical data stored across hospitals and institutions [28, 29]. At the same time, regulatory frameworks such as GDPR [11], HIPAA [35], APPI [14], and related medical-data governance initiatives [15, 10] increasingly restrict direct data centralization and cross-institutional sharing. These constraints make collaborative learning attractive, but they also require architectures that preserve data locality and limit information exposure.
Federated learning (FL) addresses part of this requirement by allowing institutions to train joint models without transferring raw data [23, 16]. However, keeping raw inputs local does not by itself eliminate privacy risk, because gradients, intermediate representations, and model updates can still leak sensitive information through inversion and inference attacks [40, 30]. Secure aggregation based on multi-party computation (MPC) mitigates this problem by revealing only aggregated client contributions [4]. In the secret-sharing-based setting considered here, this protection is information-theoretic and therefore remains secure against computationally unrestricted adversaries, including quantum ones.
MPC strengthens privacy protection but introduces a systems-level trade-off, because protecting richer intermediate representations generally increases communication and execution cost. This trade-off becomes more restrictive when a quantum post-processing module is introduced, since the quantum branch must operate on whatever protected latent representation remains practical after secure aggregation. Meanwhile, prior work on distributed and hybrid quantum learning has studied distributed quantum models [26, 31, 2, 12], tensor-network/quantum hybrid architectures [6, 21, 34, 7, 20, 32], and privacy-preserving quantum federated protocols [19]. However, these lines of work have typically focused on protocol design or model design separately, and they do not sufficiently characterize how post-aggregation quantum refinement behaves when jointly constrained by secure aggregation and representation compression.
A practical challenge is that current quantum processors provide only a limited number of usable qubits, making direct quantum processing of high-dimensional medical images unrealistic. A useful hybrid design must therefore compress the input before quantum processing while preserving task-relevant structure. Tensor networks (TNs) are a natural candidate for this role. Prior studies have established TNs as compact and structured learning models across supervised learning, generative modeling, and distributed settings [17, 25, 8, 9, 24, 1, 39, 38, 3]. In the present setting, TN compression plays a dual role: it makes small-qubit quantum processing feasible on compressed latent features and reduces the communication overhead of MPC applied to those features.
@@
Despite these promising ingredients, prior work has not sufficiently integrated representation compression, secure aggregation, and post-aggregation quantum refinement within a single privacy-aware federated pipeline. Two questions are central here: whether compressed latent representations can simultaneously improve MPC practicality and support effective small-qubit quantum processing, and whether the benefit of post-aggregation quantum refinement depends on the tensor-network frontend that produces the latent representation.
@@
To address these issues, we propose an end-to-end privacy-aware federated learning framework for medical image classification that combines client-side tensor-network encoding, MPC-secured aggregation of latent representations, and post-aggregation quantum refinement by a Quantum-Enhanced Processor (QEP). We compare three tensor-network frontends: Matrix Product State (MPS), Tree Tensor Network (TTN), and Multi-scale Entanglement Renormalization Ansatz (MERA), and evaluate how the effect of the QEP depends on the latent structure supplied by the frontend. The resulting design is intended to support small-qubit quantum processing while reducing the communication overhead associated with secure aggregation.
@@
The main contributions are threefold: (1) an end-to-end privacy-aware hybrid framework combining tensor-network latent encoding, MPC-secured aggregation, and post-aggregation quantum refinement for federated medical image classification; (2) an empirical demonstration that tensor-network compression supports both small-qubit quantum processing and lower MPC communication cost; and (3) evidence that the effect of the QEP is architecture-dependent, with TTN+QEP providing the most favorable overall profile in the present setting.
@@ II Methodology
Figure 1: Proposed TN+MPC+QEP pipeline. Client-side tensor-network encoders produce latent features, MPC-secured aggregation forms a protected global representation, and the QEP performs post-aggregation refinement before classification.
Fig. 1 shows the proposed privacy-aware federated learning framework, which consists of client-side tensor-network encoding, MPC-secured aggregation of latent representations, and post-aggregation refinement by a Quantum-Enhanced Processor (QEP). Each client maps its local input to a compact latent representation; these latent features are securely aggregated, and the resulting global representation is refined by the QEP before final classification.
@@ II-A Client-Side Tensor-Network Embedding
For the TTN and MERA branches, the input is first reorganized into non-overlapping image patches. In particular, the $28\times 28$ image is partitioned into 1616 patches of size $7\times 7$, and each patch is processed by a shared patch-wise stem network:
@@
For all frontend types, the tensor-network block maps the input to a compact latent feature through a frontend-specific structured contraction. Complex internal states are used where appropriate, while QR-based projection and intermediate normalization are introduced to improve numerical stability. The final tensor-network state is converted to a real-valued vector and projected to the shared latent dimension $d$.
@@ II-A1 MPS
In the implementation, the site tensors are parameterized in a QR-projected approximately left-canonical form,
$(A^{(k)})^{\dagger}A^{(k)}\approx I,$ (10)
and intermediate states are normalized during sequential contraction for numerical stability.
@@ II-A2 TTN
where $\cdot\|\cdot$] denotes vector concatenation. In the implementation, $\mathcal{W}^{(\ell)}$ is parameterized through QR-based projection so that
$(\mathcal{W}^{(\ell)})^{\dagger}\mathcal{W}^{(\ell)}\approx I,$ (15)
and the hidden state is normalized after each aggregation level.
@@ II-A3 MERA
followed by a split back into two updated local states. In the implementation, disentanglers are applied in alternating even and odd neighboring pairs at each level.
@@
where $\hat{h}^{(\ell)}$ denotes the disentangled local states. Both the disentangling and coarse-graining maps are implemented through QR-projected complex transformations, and hidden states are normalized after each coarse-graining step.
@@ II-B1 System Model and Threat Model
We consider $n$ clients $\{C_{i}\}_{i=1}^{n}$ and an outsourced aggregation server implemented by three non-colluding computation nodes $S_{0},S_{1},S_{2}$. Each client holds a private dataset $\mathcal{D}_{i}$ and computes a local representation $f_{i}\in\mathbb{R}^{d}$; for the MPC benchmark, a client-specific scalar coefficient $w_{i}\in\mathbb{R}$ is additionally introduced to model protected weighted aggregation and normalization. We consider both passive and active security, corresponding to Scenarios 1–3 and Scenarios 4–6, respectively.
@@ II-B2 Cost Model
In the communication model used in this paper, one secure multiplication costs $3k$ bits, one truncation costs $6k$ bits, and a full fixed-point multiplication therefore costs $9k$ bits. Secure division is modeled with communication cost $3k(k+4\theta+2)$, where $\theta$ denotes the number of refinement iterations. Active security is modeled by doubling the communication cost. Because the focus here is communication benchmarking rather than protocol derivation, we use these established costs directly.
@@ II-C Secure Aggregation via MPC
The framework includes an MPC-secured aggregation stage between client-side tensor-network embedding and post-aggregation quantum refinement. For inference evaluation, this stage is studied as part of the end-to-end architecture in Fig. 1; for communication-cost analysis, it is abstracted as a protected weighted aggregation model.
@@ II-D Quantum-Enhanced Processor
The Quantum-Enhanced Processor (QEP) refines the aggregated latent representation through quantum-state embedding, observable-based readout, and classical fusion. Let the aggregated server-side feature be denoted by
@@
In the present implementation, the quantum circuit is evaluated by statevector simulation and used as a fixed nonlinear feature transformation rather than as a trainable variational component. Accordingly, only the surrounding classical modules, including the encoder, decoder, bypass pathway, and fusion layers, are optimized during training.
From a representational viewpoint, the QEP can be interpreted as a feature-expansion mechanism that embeds a compact latent vector into a quantum Hilbert space and extracts observable-based statistics. In this setting, the effective dimensionality of the quantum feature space is determined by the number of observables used for measurement rather than by the number of circuit parameters. To ensure sufficient expressivity, it is therefore desirable that the number of quantum features $d_{q}$ be comparable to the latent dimension $d$. When using both one-body and two-body Pauli observables, this leads to a scaling relation $d_{q}=\mathcal{O}(N_{q}^{2})$, suggesting that the number of qubits should scale approximately as $N_{q}\sim\sqrt{d}$. This consideration motivates the qubit counts used in the present study.
@@ II-D1 Quantum Encoding and Observable Readout
where $d_{q}$ is determined by the number and type of observables. This yields observable-based quantum statistics from a compact latent representation without requiring a large quantum register.
@@ III Experimental Setup
We evaluated the proposed framework on PneumoniaMNIST, a binary medical image classification benchmark from the MedMNIST collection [36]. The task is to discriminate Normal from Pneumonia using resized $28\times 28$ grayscale chest X-ray images normalized to $[0,1]$. For the MPS branch, each image was flattened into a 784-dimensional vector. For the TTN and MERA branches, the same image was additionally reorganized into 1616 non-overlapping patches of size $7\times 7$, followed by a shared patch-wise stem transformation. The standard training, validation, and test splits provided by the dataset interface were used throughout.
Figure 3: Example samples from PneumoniaMNIST. The top row shows Normal cases and the bottom row shows Pneumonia cases.
To reflect the distributed setting, each split was partitioned across client branches using label-stratified client assignment. Unless otherwise noted, the number of client branches was fixed at 16. Training used a per-client local batch size of 4, yielding up to 64 samples per federated training step when all client branches contributed non-empty batches.
The overall model consisted of client-side tensor-network embedding, server-side feature gating, and an optional Quantum-Enhanced Processor (QEP), followed by a classical classifier. We considered three tensor-network frontends: Matrix Product State (MPS), Tree Tensor Network (TTN), and Multi-scale Entanglement Renormalization Ansatz (MERA). In all cases, the shared latent feature dimension was fixed to 64. The comparison is intended to evaluate practically instantiated frontend families under a common latent dimension, optimization protocol, and evaluation setting, rather than a strictly topology-isolated control.
@@
For the QEP, the quantum branch used 16 qubits, circuit depth 2, angle scale 0.5, and refresh interval 1. Quantum expectation values were computed with Qiskit Aer in statevector mode. The circuit used repeated layers of single-qubit $R_{y}$ and $R_{z}$ rotations followed by nearest-neighbor CNOT entangling gates, and the observable set included one-body Pauli terms together with selected two-body correlations. The resulting quantum statistics were classically decoded into the shared latent feature space and adaptively fused with the classical pathway.
We report both standard-threshold and threshold-optimized evaluation. Performance was assessed using Accuracy, Precision, Recall, and F1-score, and quantum runs additionally tracked internal QEP diagnostics including the mean fusion coefficient $\alpha$ and the standard deviation of the quantum-enhanced branch representation.
We also examined QEP behavior under qubit scaling and noise variation. Qubit counts were varied from 4 to 16, and noise sensitivity was evaluated at 8 qubits under noiseless, depolarizing, thermal, and mixed-noise conditions. These analyses should be interpreted as controlled studies of operating range and robustness rather than hardware-performance claims.
Based on this common setting, we define two complementary experiments: one for end-to-end predictive analysis and one for MPC communication-cost benchmarking. Within the end-to-end analysis, additional qubit-scaling and noise studies focus on the TTN-based model because the frontend comparison identifies TTN+QEP as the most balanced configuration in the present setting. This choice is intended to characterize the operating range of the quantum branch without implying that the same range transfers unchanged to other frontend types.